OpenStudy (svnnynight9):

(Algebra 2)(Quadratics with complex numbers) Which of the following expression is equal to 2x^2+8. please explain with steps I don't know how to get the answer.

OpenStudy (sshayer):

\[2x^2+8=2\left\{ x^2-\left(- 4 \right) \right\}=2\left\{ x^2-\left( 4 \times-1 \right) \right\}=2\left( x^2-4 \iota^2 \right)\] \[=2\left\{ x^2-\left( 2 \iota \right)^2 \right\}=2\left( x+2 \iota \right)\left( x-2 \iota \right)\]

OpenStudy (svnnynight9):

A)(2x+4i)(x+2i) B)(2x-4i)(x+2i) C)(2x-2i)(x+6i) D)(2x-4i)(x-2i)

OpenStudy (fwizbang):

What is 2(x-2i)?